OpenStudy (canada907cat):

Find the tangent line to y^2=x at the point (4,-2). I know that dy/dx= 1/(2y) and I know how to write the tangent line function. But since there is no x value to find the slope, would the function not exists?

OpenStudy (davejavous):

Use implicit differentiation on y^2=x

OpenStudy (canada907cat):

i did use implicit differentiation. That's how I got 1/(2y)

OpenStudy (danjs):

you can re-write the equation for y \[\large y = x ^{1/2}\] \[\frac{ dy }{ dx }=\frac{ 1 }{ 2\sqrt{x} }\]

OpenStudy (canada907cat):

okay so how would I solve it from there in terms of putting it into the tangent function?

OpenStudy (danjs):

\[y = \pm \sqrt{x}\]

OpenStudy (danjs):

the point (x , y) = (4 , -2) is given to use The slope of a tangent here at x=4 is \[\frac{ dy }{ dx }=\frac{ 1 }{ 2\sqrt{4} }=\frac{ 1 }{ 4 }\]

OpenStudy (canada907cat):

okay so I would put 1/4 into the tangent formula correct?

OpenStudy (danjs):

sorry you need the negative slope for that point \[\frac{ dy }{ dx }=-1/4\]

OpenStudy (danjs):

When x=4, y=-2, that is the negative root

OpenStudy (canada907cat):

Okay so would I do anything with 1/(2y) or would I just put in the -1/4 for the slope in the tangent function. Oh okay.

OpenStudy (danjs):

yeah so you have the slope and a point the line goes through.. line slope = -1/4 through point (4 , -2) y - (-2) = (-1/4)*(x - 4)

OpenStudy (canada907cat):

Don't you mean T(x)= -2-1/4(x-4)?

OpenStudy (danjs):

you can clean it up a bit.. y - (-2) = (-1/4)*(x - 4) \[y = \frac{ -1 }{ 4 }x-1\]